A 50.0-cm-long hollow nichrome tube of inner diameter 2.60mm , outer diameter 5.80mm is connected to a 4.50V battery. What is the current in the tube?
Answer #1

Area = 8.446/4 x 10^-5 = 2.111 x 10^-5

So, R = 1.5 x 10^-6 * 0.5/ 2.111 x 10^-6 = 0.4263 ohm

So, i = 4.5/0.4263 = 10.555 A